OpenStudy (anonymous):

the endpoints of a segmant are (10,-5) and (3,y). What is a possible value of y if the segment length is 25?

OpenStudy (amistre64):

pythag it :)

OpenStudy (amistre64):

25^2 = (10-3)^2 + (-5-y)^2

OpenStudy (anonymous):

YOU!!! >: O

OpenStudy (amistre64):

-sqrt(225 - 49) -5 = y maybe?

OpenStudy (amistre64):

err... 625 = 25^2 right lol

OpenStudy (amistre64):

lets go with -29

OpenStudy (anonymous):

haha thanks for showing your work, that helps me a lot!!